# Completely (Quasi-)Uniform Nested Boolean Steiner Quadruple Systems

### v2

@@ Abstract
The notion of completely uniform pairings is further generalized for $t$-designs with $t\geq 2$. As an application, completely uniform nested $2$-$(2^{m},4,3)$ designs give rise to fractional repetition codes with zero skip cost, requiring fewer storage nodes than constructions based on SQSs. In addition, small examples are provided for non-Boolean orders, establishing the existence of completely uniform nested SQS(v)(v) for all $v\leq 50$.
@@ 1 Introduction
From the viewpoint of design theory, direct constructions for SQSs (and, more generally, for $t$-designs with $t\geq 3$) are rather limited. Among known families, Boolean SQSs are classical and have rich algebraic and finite-geometric structure, so they provide a natural entry point to completely uniform pairings, for which no direct algebraic or finite-geometric constructions were previously known. This work can be regarded as a first step in this direction and may help address other open problems raised in Chee et al. [8].
@@ 2.3 Necessary conditions for the existence of completely (quasi-)uniform nested SQSs
While the concept of quasi-uniform nested SQSs was introduced in [8], their constructions and further properties were not extensively examined. As an analogue of Theorem 2.5, we derive below a necessary condition for the existence of a completely quasi-uniform nested SQS.
@@ 2.4 Generalization of completely (quasi-)uniform nested designs
It is worth noting that the case $k=4$, $\lambda=3$, and $\ell=2$ is also closely related to the classical problem of whist tournaments, whose study dates back to Moore’s work in 1896 [30, 31]. More precisely, a whist tournament of order $v=4n$ is exactly a completely uniform resolvable $2$-nested 22-$(v,4,3)$ design (see [2, 32] for details).
@@ 3 Construction for completely (quasi-)uniform nested $\operatorname{SQS}(2^{m})$
Fractional repetition (FR) codes were introduced by El Rouayheb and Ramchandran [11] as a combinatorial tool for distributed storage systems. They are used as the inner component of DRESS (Distributed Replication based Exact Simple Storage) codes [34], which combine an outer MDS code with an inner FR code. In an FR code, symbols are replicated across storage nodes in a structured manner. A failed node can be repaired by transferring symbols from other surviving nodes, called helper nodes, without additional computation. This structure enables simple and efficient repair in large-scale storage systems, which also balances the I/O and bandwidth load among helper nodes.
@@ 6 Concluding remarks
This work proposes explicit constructions for completely (quasi-)uniform nested SQSs based on Boolean $\operatorname{SQS}(2^{m})$, resolving two open problems from [8] by constructing infinite families of such designs for odd $m\geq 3$ (completely uniform) and even $m\geq 4$ (completely quasi-uniform).
The constructions presented in this work rely heavily on the rich algebraic and geometric structure of the Boolean SQSs, in particular their affine invariance ($2$-transitivity), rotational automorphisms, and resolvable properties.
@@ Acknowledgements
The authors would like to thank the anonymous reviewers for their constructive comments, which clarified the significance of this work and improved the overall structure of the paper.
This work was partly supported by JSPS KAKENHI Grant Nos. 22K13949, 22K11943, and 24K14819.

### v3

@@ 1 Introduction
From the viewpoint of design theory, direct constructions for SQSs (and, more generally, for $t$-designs with $t\geq 3$) are rather limited. Among known families, Boolean SQSs are classical and have rich algebraic and finite-geometric structure, so they provide a natural entry point to completely uniform pairings, for which no direct algebraic or finite-geometric constructions were previously known. This study can be regarded as a first step in this direction and may help address other open problems raised in Chee et al. [8].
@@ 2.3 Necessary conditions for the existence of completely (quasi-)uniform nested SQSs
While the concept of quasi-uniform nested SQSs was introduced in [8], their constructions and further properties were not extensively examined. As an analog of Theorem 2.5, we derive below a necessary condition for the existence of a completely quasi-uniform nested SQS.
@@ 2.4 Generalization of completely (quasi-)uniform nested designs
It is worth noting that the case $k=4$, $\lambda=3$, and $\ell=2$ is also closely related to the classical problem of whist tournaments, whose study dates back to Moore’s study in 1896 [30, 31]. More precisely, a whist tournament of order $v=4n$ is exactly a completely uniform resolvable $2$-nested 22-$(v,4,3)$ design (see [2, 32] for details).
@@ 3 Construction for completely (quasi-)uniform nested $\operatorname{SQS}(2^{m})$
FR codes were introduced by El Rouayheb and Ramchandran [11] as a combinatorial tool for distributed storage systems. They are used as the inner component of DRESS (Distributed Replication based Exact Simple Storage) codes [34], which combine an outer MDS code with an inner FR code. In an FR code, symbols are replicated across storage nodes in a structured manner. A failed node can be repaired by transferring symbols from other surviving nodes, called helper nodes, without additional computation. This structure enables simple and efficient repair in large-scale storage systems, which also balances the I/O and bandwidth load among helper nodes.
@@ 6 Concluding remarks
This study proposes explicit constructions for completely (quasi-)uniform nested SQSs based on Boolean $\operatorname{SQS}(2^{m})$, resolving two open problems from [8] by constructing infinite families of such designs for odd $m\geq 3$ (completely uniform) and even $m\geq 4$ (completely quasi-uniform).
The constructions presented in this study rely heavily on the rich algebraic and geometric structure of the Boolean SQSs, in particular their affine invariance ($2$-transitivity), rotational automorphisms, and resolvable properties.
@@ Acknowledgements
The authors would like to thank the anonymous reviewers for their constructive comments, which clarified the significance of this study and improved the overall structure of the paper.
This study was partly supported by JSPS KAKENHI Grant Nos. 22K13949, 22K11943, and 24K14819.